Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $51,000 $116,000
Machine hours 1,410 1,330
Setup hours 96 400


Total estimated overhead costs are $304,000. Overhead cost allocated to the machining activity cost pool is $195,000, and $109,000 is allocated to the machine setup activity cost pool.

Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.)

Predetermined overhead rate

enter the predetermined overhead rate in percentages of direct labor cost rounded to 2 decimal places

% of direct labor cost

eTextbook and Media

  

  

Compute the overhead rates using the activity-based costing approach. (Round answers to 2 decimal places, e.g. 12.25.)

Machining

$enter a dollar amount per machine hour rounded to 2 decimal places

per machine hour
Machine setup

$enter a dollar amount per setup hour rounded to 2 decimal places

per setup hour

eTextbook and Media

  

  

Determine the difference in allocation between the two approaches. (Round answers to 0 decimal places, e.g. 1,225.)

Traditional costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

Activity-based costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Saddle
Answer 1
Calculation of Predetermined OH Rate Standard Custom Total Note
Direct Labor cost                      51,000.00 116,000.00    167,000.00 A
Total Manufacturing overhead    304,000.00 B
Predetermined OH Rate 182.04% C=B/A
Total Allocated OH                      92,840.00 211,166.00    304,006.00 D=C*A
Answer 2 E F G H=F+G I=H/E
Calculation of ABC rates for overhead Activity usage
Activity cost pool Activity Measure Cost Standard Custom Total ABC rates
Machining activity Machine hours 195,000.00         1,410.00            1,330.00         2,740.00               71.17
Machine Setups Number of setups 109,000.00              96.00               400.00            496.00             219.76
Total Cost assigned 304,000.00
Allocation table
Standard See I See F J=F*I
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned
Machine Setups Number of setups             71.17         1,410.00       100,350.00
Mixtures Number of mixtures           219.76              96.00         21,097.00
Total Cost assigned       121,447.00
Deluxe See I See G K=G*I
Activity cost pool Activity Measure ABC Rates Activity used Cost assigned
Machine Setups Number of setups             71.17         1,330.00         94,656.00
Mixtures Number of mixtures           219.76            400.00         87,904.00
Total Cost assigned       182,560.00
Answer 3 Standard Deluxe Note
Overhead assigned- Traditional costing                      92,840.00 211,166.00 See D
Overhead assigned- ABC                    121,447.00 182,560.00 L= J,K
Difference                    (28,607.00)     28,606.00 M=D-L